In a frequency distribution, the first class interval begins at 18. The midpoint of the first class interval is 19.5, and the midpoint of the last class interval is 49.5. How many class intervals are there? A. 11 B. 17 C. 22 D. 33 E. 34
Answer #1

mid point = (lower limit + upper limit )/2 = 19.5

(18+ upper limit)/2 = 19.5

upper limit = 21

so, first class interval is (18-21)

so, class width =21-18=3

so, other mid point will be

19.5
22.5
25.5
28.5
31.5
34.5
37.5
40.5
43.5
46.5
49.5
hence, total intervals = 11
Answer #2
Suppose the instructor announced that your class of 65 students has a mean grade of 73.2 this implies that the sum of all the 65 grades is 4758 is this true or false
source: Statistics exam
answered by: Rasha